OpenStudy (anonymous):

what is the minimum value of the function f(x)=2x^2-4x+1?

OpenStudy (goformit100):

f(x)=2x^2-4x+1 differentiate it.

OpenStudy (goformit100):

After then what to do will be said by @satellite73 SIR

OpenStudy (anonymous):

or look for the second coordinate of the vertex first coordinate is always \(-\frac{b}{2a}\)

OpenStudy (goformit100):

@satellite can you Help in my preparation in My Exam ?

OpenStudy (anonymous):

WHAT DOES MINIMUM VALUE MEAN?

OpenStudy (anonymous):

What does minimum value mean?

OpenStudy (anonymous):

it means the smallest value that \(y\) can be

OpenStudy (anonymous):

so what would the minumum value in this eqution be

OpenStudy (anonymous):

\[f(x)=2x^2-4x+1 \] \[f(x)=ax^2+bx+c\] in this case \(a=2,b=-4\) so \(\frac{-b}{2a}=-\frac{-4}{2\times 2}=1\) that is the first coordinate of the vertex second is \(f(1)\)

OpenStudy (anonymous):

 hihihi hope this helps